For a record playing at 78 rpm, find the frequency and the angular velocity.

frequency =

angular velocity=
Answer #1

N = 78 rpm

frequency = N/60

= 78/60 = 1.3 Hz

angular velocity = 2pi*f

= 2pi*1.3 = 8.168 rad/s